# Noah Porcelli

### v1

@@ 3 Moduli spaces
In this section, we construct moduli spaces $\mathcal{M}$ and $\mathcal{P}_{r}$ for $r\geq 0$, similarly to [10, Sections 4.1 & 5.3] and [6].
Let $D$ be a smooth convex subdomain of $\mathbb{C}$, lying inside $\mathbb{R}+i[0,1]$ and containing $[-1,1]+i[0,1]$.
@@
We define $Z_{-}\mathrel{\mathop{:}}=D\cup\left([0,\infty)+i[0,1]\right)$.
@@
$Z_{r}\mathrel{\mathop{:}}=\left(D-r\right)\cup\left([-r,r]+i[0,1]\right)\cup\left(D+r\right)$
@@
Recall the (topological) energy of a $J$-holomorphic map $u\mathrel{\mathop{:}}\Sigma\to M$ from a Riemann surface $\Sigma$ is the quantity
$E(u)\mathrel{\mathop{:}}=\int_{\Sigma}u^{*}\omega$
@@
For $r\geq 0$, we define $\mathcal{P}_{r}$ to be the space of $J$-holomorphic maps $u\mathrel{\mathop{:}}Z_{r}\to M$ with finite energy, with moving Lagrangian boundary conditions:
@@
We define $\mathcal{M}$ to be the space of $J$-holomorphic maps $u\mathrel{\mathop{:}}Z_{-}\to M$ which have finite energy, satisfying (2) (with $Z_{r}$ replaced by $Z_{-}$).
@@
Note that we do not quotient by any automorphisms in the definitions of these moduli spaces; this does not interfere with compactness, by the argument in [10, Lemma 4.5].
[10, Lemma 4.3] shows there is an a priori uniform upper energy bound on the energy of any $u\in\mathcal{M}$ or $u\in\mathcal{P}_{r}$, independent of $r$ (dependent however on the choice of $\{\psi^{t}\}_{t}$).
Using relative exactness, Gromov compactness [9, Theorem 4.6.1] implies that
@@
Let $C\geq 0$ and $x\in[-C,C]$, and let $ev_{x}\mathrel{\mathop{:}}\mathcal{P}_{C}\to L$ send $u$ to $u(x)$. Then under the orientation conditions of Theorem 1.5 for $R^{*}=H^{*}(\cdot;\mathbb{Z}/2),H^{*}(\cdot;\mathbb{Z})$ or $K^{*}$, the induced map $ev^{*}_{x}\mathrel{\mathop{:}}R^{*}(L)\to R^{*}(\mathcal{P}_{C})$ is injective.
This is proved in [6, Theorem 3.1] in the case $R^{*}$ is (a Čech model for) $H^{*}(\cdot;\mathbb{Z}/2)$ and [5, Theorem 2.12] in the other cases. Since $\mathcal{P}_{C}$ can in generally be extremely singular (and in particular, not necessarily homotopy equivalent to a CW complex), we must be careful with exactly what model of $R^{*}(\mathcal{P}_{C})$ we use; see [5, Section 2.1] for a discussion about this point. However all reasonable models for $R^{*}(\cdot)$ agree when applied to $L$, since $L$ admits the structure of a finite CW complex, so this does not affect how we apply Proposition 3.4.
@@ 4 Pinching
$d_{u}(x)\mathrel{\mathop{:}}=\sup_{t\in[0,1]}d\left(u(x),u(x+ti)\right)$
@@
Since we assumed $L$ and $L^{\prime}$ intersect transversally, for fixed $v\in\mathcal{M}$ this holds for $\lambda$ sufficiently large, by the standard exponentian decay estimate for holomorphic curves on their strip-like ends; see e.g. [11]; we may choose $\lambda>0$ large enough that this holds for all $v\in\mathcal{M}$ since $\mathcal{M}$ is compact. ∎
@@
The right hand side of (5) is $\leq\varepsilon/2$ by Lemma 4.2, whereas the right hand side is $>\varepsilon$ for all $n$, by (3), giving a contradiction. ∎
@@
Let $C$ and $x_{0}$ be as in Proposition 4.1 and let $ev_{x_{0}},ev_{x_{0}+i}\mathrel{\mathop{:}}\mathcal{P}_{C}\to L$ be the maps which evaluate at $x_{0}$ and $x_{0}+i$ respectively
@@
Proof.
We write down an explicit homotopy; this can be thought of as “evaluating around the boundary of $Z_{C}$ from $x_{0}$ to $x_{0}+i$”. Let $\gamma\mathrel{\mathop{:}}[0,1]\to\partial Z_{C}$ be any path from $x_{0}$ to $x_{0}+i$, and let $r_{t}=\operatorname{Im}\gamma(t)$ for $t\in[0,1]$. Then we define a homotopy $H\mathrel{\mathop{:}}[0,1]\times\mathcal{P}_{C}\to L^{\prime}$ to send $(t,u)$ to
@@
Proof.
By Proposition 4.1 and (1), for all $u\in\mathcal{P}_{C}$, $\pi\circ ev_{x_{0}+i}(u)$ and $ev_{x_{0}}(u)$ are of distance at most $2\varepsilon$ from each other. Since $2\varepsilon$ is less than the injectivity radius, the following diagram commutes, via a homotopy which follows the unique small geodesic between the two endpoints:
[…]
Combined with (6) and Lemma 2.1, this implies the result. ∎
@@
Let $R$ be the cohomology theory $H^{*}(\cdot;\mathbb{Z}/2),H^{*}(\cdot;\mathbb{Z})$ or $K^{*}$, and assume the relevant orientation condition from Theorem 1.7 holds.
We let $C$ and $x_{0}$ be as in Proposition 4.1. We define a family of maps $\{\phi^{t}\}_{t\in[0,1]}$ to be given by $\phi^{t}=\psi^{2t}$ for $0\leq t\leq\frac{1}{2}$ and to follow the unique small geodesic between $\psi^{1}(x)$ and $\phi(x)$ for $x\in M$ and $\frac{1}{2}\leq t\leq 1$. Then $\phi^{1}=\phi$ and $\phi^{0}$ is the identity.
Consider the sweep-out map $S\mathrel{\mathop{:}}L\times[0,1]\to M$, sending $(x,t)$ to $\phi^{t}(x)$. This sends $L\times\{0,1\}$ to $L$ and therefore a map of pairs $(L\times[0,1],L\times\{0,1\})\to(M,L)$; applying $R^{*}$ and applying the suspension isomorphism gives a map
$\sigma\mathrel{\mathop{:}}R^{*+1}(M,L)\to R^{*}(L)$
Let $\partial\mathrel{\mathop{:}}R^{*}(L)\to R^{*+1}(M,L)$ be the boundary map.
@@
$\phi^{*},(Id+\partial\circ\sigma)\mathrel{\mathop{:}}R^{*}(M,L)\to R^{*}(M,L)$
@@
Let $\overline{M}=M\cup_{L\times\{0\}}L\times[0,1]$ be the mapping cylinder of the inclusion $L\hookrightarrow M$. Then $\phi$ induces a map of pairs $(\overline{M},L\times\{1\})$ to itself, and the action of this map on $R^{*}$ is the same as that of the action of $\phi$ on $(M,L)$.
@@
Define $H\mathrel{\mathop{:}}\overline{M}\times[0,1]_{s}\to\overline{M}$ to send $(x,s)\in M\times[0,1]$ to $\phi^{s}(x)$ and $(x,t,s)\in L\times[0,1]^{2}$ to
@@
We a map define a map $T\mathrel{\mathop{:}}\mathcal{P}_{C}\times[0,1]\to M$ as follows. Let $\gamma\mathrel{\mathop{:}}[0,1]\to\partial Z_{C}$ be a path from $x_{0}$ to $x_{0}+i$ and let $r_{t}=\operatorname{Im}\gamma(t)$. Then we define $T(u,t)$ to be $u\circ\gamma(2t)$ for $0\leq t\leq\frac{1}{2}$ and to follow the unique small geodesic between $ev_{x_{0}+i}(u)$ and $\phi\circ(\psi^{1})^{-1}\circ ev_{x_{0}+i}(u)$ for $\frac{1}{2}\leq t\leq 1$.
We further define a map $F\mathrel{\mathop{:}}\mathcal{P}_{C}\times[0,1]\to L\times[0,1]$ by
$F(u,t)\mathrel{\mathop{:}}=\begin{cases}\left(\left(\psi^{r_{2t}}\right)^{-1}\circ u\circ\gamma(2t),t\right)&\textrm{ if }0\leq t\leq\frac{1}{2}\\ \left(\left(\psi^{1}\right)^{-1}\circ ev_{x_{0}+i}(u),t\right)&\textrm{ if }\frac{1}{2}\leq t\leq 1\end{cases}$
@@
Proof.
We define a map $S^{\prime}\mathrel{\mathop{:}}L\times[0,1]\to M$ to send $(x,t)$ to $\psi^{r_{2t}}(x)$ if $0\leq t\leq\frac{1}{2}$ and to follow the unique small geodesic between $\psi^{1}(x)$ and $\phi(x)$ for $\frac{1}{2}\leq t\leq 1$. By construction, the following diagram commutes:
[…]
$S^{\prime}$ and $S$ are homotopic (relative to $L\times\{0,1\}$) since they are the same up to a reparametrisation of $[0,\frac{1}{2}]$.
@@
Follows from Lemmas 6.1 and 6.3. ∎

### v2

@@ 4 Moduli spaces
In this section, we construct moduli spaces $\mathcal{M}$ and $\mathcal{P}_{r}$ for $r\geq 0$, similarly to [17, Sections 4.1 & 5.3] and [9].
[…]
@@
We define $Z_{-}\mathrel{\mathop{\ordinarycolon}}=D\cup\left([0,\infty)+i[0,1]\right)$.
@@
$Z_{r}\mathrel{\mathop{\ordinarycolon}}=\left(D-r\right)\cup\left([-r,r]+i[0,1]\right)\cup\left(D+r\right)$
@@
A compatible almost complex structure on $(M,\omega)$ is a vector bundle automorphism $J$ of the tangent bundle $TM\to M$, such that $J^{2}=-\mathrm{Id}_{TM}$, and that the tensor $\omega(\cdot,J\cdot)$ defines a $J$-invariant Riemannian metric on $M$.
It is standard that compatible almost complex structures always exist, and so we may choose such a $J$ on $M$, which is additionally convex at infinity (a “niceness at infinity”-type condition, see [22, Section 7(b)]) if $M$ is Liouville. We fix this $J$ throughout the rest of the paper.
Recall a smooth map $u\mathrel{\mathop{\ordinarycolon}}\Sigma\to M$ from a Riemann surface $(\Sigma,j)$ with boundary is $J$-holomorphic if $du\circ j=J\circ du$. The (topological) energy of such a $J$-holomorphic $u\mathrel{\mathop{\ordinarycolon}}\Sigma\to M$ is the quantity
$E(u)\mathrel{\mathop{\ordinarycolon}}=\int_{\Sigma}u^{*}\omega$
We take $\psi$ to be as defined in Section 3. Since $\psi$ is a Hamiltonian diffeomorphism, we may choose a Hamiltonian isotopy $\{\psi^{t}\}_{t\in[0,1]}$, with $\psi^{0}=\mathrm{Id}_{M}$ and $\psi^{1}=\psi$. We keep this fixed throughout the rest of the paper.
Definition 4.3.
For $r\geq 0$, we define $\mathcal{P}_{r}$ to be the space of $J$-holomorphic maps $u\mathrel{\mathop{\ordinarycolon}}Z_{r}\to M$ with finite energy, with moving Lagrangian boundary conditions:
@@
We define $\mathcal{M}$ to be the space of $J$-holomorphic maps $u\mathrel{\mathop{\ordinarycolon}}Z_{-}\to M$ which have finite energy, satisfying (3) (with $Z_{r}$ replaced by $Z_{-}$). We equip $\mathcal{M}$ with the weak $C^{\infty}$-Whitney topology.
@@
Note that we do not quotient by any automorphisms in the definitions of these moduli spaces; this does not interfere with compactness, by the argument in [17, Lemma 4.5].
[17, Lemma 4.3] shows there is an a priori uniform upper energy bound on the energy of any $u\in\mathcal{M}$ or $u\in\mathcal{P}_{r}$, independent of $r$ (dependent however on the choice of $\{\psi^{t}\}_{t}$).
@@
Proof.
Compactness of each $\mathcal{P}_{r}$ is proved in [17, Lemma 4.5] (where $\mathcal{P}_{r}$ is denoted $\mathcal{U}_{r}$); the same argument applies also to $\mathcal{M}$. Note that this requires both relative exactness of $L$ as well as cylindricity at infinity of $J$.
The final claim is an application of [15, Theorem 4.6.1]. Explicitly in the notation of loc. cit., we set $\nu=n$, $\Sigma=Z_{-}$, $J^{n}=J$, $\Omega^{n}=(Z_{r_{n}}\cap\mathbb{C}_{\mathrm{Re}<0})+n\subseteq\Sigma$ and $u^{n}=u_{n}(\cdot+r_{n})$. Finally relative exactness of $L$ allows us to conclude the set $Z$ arising in the conclusion of loc. cit. must be empty (since no bubbling can occur). ∎
@@
Let $C\geq 0$ and $x\in[-C,C]$, and let $ev_{x}\mathrel{\mathop{\ordinarycolon}}\mathcal{P}_{C}\to L$ send $u$ to $u(x)$. Then under the orientation conditions of Theorem 1.11 for $R^{*}=H^{*}(\cdot;\mathbb{Z}/2),H^{*}(\cdot;\mathbb{Z})$ or $K^{*}$, the induced map $ev^{*}_{x}\mathrel{\mathop{\ordinarycolon}}R^{*}(L)\to R^{*}(\mathcal{P}_{C})$ is injective.
Proof.
This is proved in [9, Theorem 3.1] in the case $R^{*}$ is (a Čech model for) $H^{*}(\cdot;\mathbb{Z}/2)$ and [8, Theorem 2.12] in the other cases. ∎
Remark 4.6.
Since $\mathcal{P}_{C}$ can in general be extremely singular (and in particular, not necessarily homotopy equivalent to a CW complex), we must be careful with exactly what model of $R^{*}(\mathcal{P}_{C})$ we use; see [8, Section 2.1] for a discussion about this point. However all reasonable models for $R^{*}(\cdot)$ agree when applied to $L$, since $L$ admits the structure of a finite CW complex, so this does not affect how we apply Proposition 4.5.
@@ 5 Pinching
$d_{u}(x)\mathrel{\mathop{\ordinarycolon}}=\sup_{t\in[0,1]}d\left(u(x),u(x+ti)\right)$
@@
Since we assumed $L$ and $L^{\prime}$ intersect transversally, for fixed $v\in\mathcal{M}$ this holds for $\lambda$ sufficiently large, by the standard exponential decay estimate for holomorphic curves on their strip-like ends, see e.g. [18]; we may choose $\lambda>0$ large enough that this holds for all $v\in\mathcal{M}$ since $\mathcal{M}$ is compact. ∎
@@
The right hand side of (6) is $\leq\varepsilon/2$ by Lemma 5.2, whereas the left hand side is $>\varepsilon$ for all $n$, by (4), giving a contradiction. ∎
@@
Let $C$ and $x_{0}$ be as in Proposition 5.1 and let $ev_{x_{0}},ev_{x_{0}+i}\mathrel{\mathop{\ordinarycolon}}\mathcal{P}_{C}\to L$ be the maps which evaluate at $x_{0}$ and $x_{0}+i$ respectively
@@
${\mathcal{P}_{C}}${${L^{\prime}}$$\scriptstyle{ev_{x_{0}}}$$\scriptstyle{ev_{x_{0}+i}}$$\scriptstyle{\psi^{1}}$ (7)
Proof.
@@
${\mathcal{P}_{C}}${${L}$$\scriptstyle{ev_{x_{0}}}$$\scriptstyle{ev_{x_{0}}}$$\scriptstyle{\phi}$ (8)
Proof.
By Proposition 5.1 and (2), for all $u\in\mathcal{P}_{C}$, $\pi\circ ev_{x_{0}+i}(u)$ and $ev_{x_{0}}(u)$ are of distance at most $2\varepsilon$ from each other. Since $2\varepsilon$ is less than the injectivity radius, the following diagram commutes, via a homotopy which follows the unique small geodesic (which lies in $L$, since $L\subseteq M$ is totally geodesic) between the two endpoints:
@@
Combined with Lemma 6.1 and Lemma 3.1, this implies the result. ∎
@@
Let $R$ be the cohomology theory $H^{*}(\cdot;\mathbb{Z}/2),H^{*}(\cdot;\mathbb{Z})$ or $K^{*}$, and assume the relevant orientation condition from Theorem 1.14 holds for $R$.
We let $C$ and $x_{0}$ be as in Proposition 5.1. We define a family of maps $\{\phi^{t}\mathrel{\mathop{\ordinarycolon}}M\to M\}_{t\in[0,1]}$ to be given by $\phi^{t}=\psi^{2t}$ for $0\leq t\leq\frac{1}{2}$, and to follow the unique small geodesic between $\psi^{1}(x)$ and $\phi(x)$ for $x\in M$ and $\frac{1}{2}\leq t\leq 1$. Then $\phi^{1}=\phi$ and $\phi^{0}$ is the identity.
[…]
$\sigma\mathrel{\mathop{\ordinarycolon}}R^{*+1}(M,L)\to R^{*}(L)$
Let :$\partial\mathrel{\mathop{\ordinarycolon}}R^{*}(L)\to R^{*+1}(M,L)$ be the boundary map.
@@
$\phi^{*},(Id+\partial\circ\sigma)\mathrel{\mathop{\ordinarycolon}}R^{*}(M,L)\to R^{*}(M,L)$
@@
[…]
@@
Define $H\mathrel{\mathop{\ordinarycolon}}\overline{M}\times[0,1]_{s}\to\overline{M}$ to send $(x,s)\in M\times[0,1]$ to $\phi^{s}(x)$ and $(x,t,s)\in L\times[0,1]^{2}$ to
@@
We define a map $T\mathrel{\mathop{\ordinarycolon}}\mathcal{P}_{C}\times[0,1]\to M$ as follows. Let $\gamma\mathrel{\mathop{\ordinarycolon}}[0,1]\to\partial Z_{C}$ be a path from $x_{0}$ to $x_{0}+i$ and let $r_{t}=\operatorname{Im}\gamma(t)$. Then we define $T(u,t)$ to be $u\circ\gamma(2t)$ for $0\leq t\leq\frac{1}{2}$ and to follow the unique small geodesic between $ev_{x_{0}+i}(u)$ and $\phi\circ(\psi^{1})^{-1}\circ ev_{x_{0}+i}(u)$ for $\frac{1}{2}\leq t\leq 1$.
We further define a map $F\mathrel{\mathop{\ordinarycolon}}\mathcal{P}_{C}\times[0,1]\to L\times[0,1]$ by
$F(u,t)\mathrel{\mathop{\ordinarycolon}}=\begin{cases}\left(\left(\psi^{r_{2t}}\right)^{-1}\circ u\circ\gamma(2t),t\right)&\textrm{ if }0\leq t\leq\frac{1}{2}\\ \left(\left(\psi^{1}\right)^{-1}\circ ev_{x_{0}+i}(u),t\right)&\textrm{ if }\frac{1}{2}\leq t\leq 1\end{cases}$
@@
${\mathcal{P}_{C}\times[0,1]}$${L\times[0,1]}$${M}$$\scriptstyle{F}$$\scriptstyle{T}$$\scriptstyle{S}$
Proof.
@@
${\mathcal{P}_{C}\times[0,1]}$${L\times[0,1]}$${M}$$\scriptstyle{F}$$\scriptstyle{T}$$\scriptstyle{S^{\prime}}$
$S^{\prime}$ and $S$ are homotopic (relative to $L\times\{0,1\}$) since they are the same up to a reparametrisation of $[0,\frac{1}{2}]$. ∎
@@
Follows from Lemmas 7.1 and 7.3. ∎